# A Two-Step Procedure to Detect Cosmological Gravitational Wave Backgrounds with Next-Generation Terrestrial Gravitational-Wave Detectors

### v1

@@ Abstract
Several methods have been proposed to bypass the CBC foreground. Given the high fraction of individually resolved CBC signals with XG detectors, one could simply fit and subtract them from the data in the frequency domain to lower the foreground Regimbau et al. (2017); Sachdev et al. (2020). However, the recovery of detected signals is never perfect, and the pile-up of residuals from imperfect subtraction can produce an effective foreground that is comparable to the original SGWB before removal Zhou et al. (2023, 2022); Song et al. (2024). Techniques have been proposed to further reduce this effective foreground in the frequency domain, from applying a projection method Cutler and Harms (2006); Sharma and Harms (2020) to estimating and subtracting the expected residual power Pan and Yang (2023). Ref. Belgacem et al. (2024) showed that subtraction results in the frequency domain are actually significantly more optimistic if one abandons the linear-signal approximation in favor of more realistic estimates of the residuals. On the other hand, Refs. Zhong et al. (2023, 2024) proposed a time-frequency approach to notch-out the detected signals by masking the associated pixels. They showed that this method could reduce the CBC foreground by one order of magnitude even when residuals from imperfect recovery are taken into account Zhong et al. (2024).
@@
Astrophysical population and setup. For both BBHs and BNSs, we assume the same population models as in Refs. Zhong et al. (2023, 2024). We employ the PowerLaw+Peak model from the latest LIGO-Virgo-KAGRA (LVK) catalog Abbott et al. (2023) to characterize the BBH mass distribution, while we assume that the BNS masses are uniformly distributed within the range $[1,2]\,\rm{M_{\odot}}$ Abbott et al. (2023); Landry and Read (2021).We assume zero spins and an isotropic distribution of the orbital orientation and sky position for all the binaries. To obtain the redshift distribution, we convolve the star formation rate (SFRFinkel et al. (2022)
@@
with a power-law time delay distribution $p(t_{d})\propto t_{d}^{p}$ with $p=-1$. The source-frame merger rate $R_{m}(z)$ then reads
@@
Let us consider a set of data segments $\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}$ from $N_{\mathrm{obs}}$ individual BNS detections and the cross-correlation spectrum $\hat{C}^{\star}(f)$ after notching out the resolvable BBH signals (see the Supplement for the definition of $\hat{C}^{\star}(f)$). Their joint likelihood can be written as Callister et al. (2020)
($\mathscr{L}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}},\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})=\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})\times\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})\,,$ (4)
where $\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})$ is the usual hierarchical likelihood for population inference on the detected events Mandel et al. (2019); Taylor and Gerosa (2018), while $\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})$ is the Gaussian likelihood for stochastic searches Mandic et al. (2012); Callister et al. (2017). Here $\bm{\Lambda}\bm{}\mathrm{CGWB}=\{\Omega_{\mathrm{ref}},\alpha\}$ encodes the parameters characterizing the CGWB, while $\bm{\Lambda}\bm{}\mathrm{BNS}$ is the set of hyperparameters defining the BNS population, which affect both the individually detected events and the SGWB from unresolved BNS signals. To simplify the inference, we assume that the BNS mass distribution is perfectly known and that the angular parameters are isotropically distributed. Hence, the only BNS hyperparameters $\bm{\Lambda}\bm{}\mathrm{BNS}$ that we are inferring are the ones characterizing the redshift distribution. In particular, we consider two cases:
@@
    A simplistic known SFR scenario where we assume that the SFR is completely known, and all that is left to determine are the local merger rate, $\mathcal{R}_{0}$, and the power-law index of the time-delay distribution, $p$. Then the joint inference involves a total of four parameters $\{\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB}\}=\{\mathcal{R}_{0},p,\Omega_{\mathrm{ref}},\alpha\}$.
@@
In vanilla searches for a single, dominant SGWB, its SNR is straightforward to compute, and it is a good figure of merit to estimate whether the target SGWB is strong enough to be detectable relative to the detector noise Abbott et al. (2023); Allen and Romano (1999); Abbott et al. (2021). In our case, instead, we are jointly fitting for the BNS hyperparameters from individually detected sources plus two distinct SGWBs, and our target is, in fact, the subdominant CGWB. Hence, we turn to a full Bayesian approach and we study the inferred posterior to find evidence for CGWB detection.
@@
Joint search: Known SFR $319^{+25}_{-24}$ $-1.00^{+0.04}_{-0.04}$ $1.92^{+0.58}_{-0.52}\times 10^{-11}$ $-0.10^{+0.32}_{-0.37}$
Joint search: Unknown SFR $305^{+48}_{-47}$ $2.24^{+0.19}_{-0.16}$ $1.43^{+0.28}_{-0.24}$ $1.64^{+0.07}_{-0.07}$ $2.04^{+0.66}_{-0.66}\times 10^{-11}$ $-0.06^{+0.31}_{-0.39}$
Ideal search - - - - - $1.95^{+0.35}_{-0.34}\times 10^{-11}$ $-0.14^{+0.29}_{-0.31}$
@@
Sensitivity of the CGWB search. In order to further assess the constraining power of our procedure, in Fig. 3 we compare the noise level for stochastic searches, the uncertainty on the BNS foreground, and the residual BBH foreground after notching out the resolved signals. The baseline noise level $\sigma_{n}(f)$ is computed through Eqs. (S1) and (S2) of the Supplement. The uncertainty on the inferred BNS foreground is estimated by computing energy spectra $\Omega_{\mathrm{BNS}}(f)$ for all different posterior samples and the resulting standard deviation $\sigma_{\Omega_{\mathrm{BNS}}}(f)$ at each frequency (see the Supplement for the description of $\Omega_{\mathrm{BNS}}(f)$ calculation). The sensitivity of stochastic searches scales as $\sim 1/\sqrt{T_{\mathrm{obs}}}$ Allen and Romano (1999), so we show the comparison for different values of $T_{\mathrm{obs}}$ by simply rescaling $\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T^{\prime}_{obs}}}(f)=\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T_{obs}}}(f)\left(\sqrt{T_{\mathrm{obs}}}/\sqrt{T^{\prime}_{\mathrm{obs}}}\right)$. We also explicitly check this scaling by running the same full simulation using fewer data, in which case we do observe the width of the posterior distributions of parameters scale as $1/\sqrt{T_{\mathrm{obs}}}$ that further makes $\sigma_{\Omega_{\mathrm{BNS}}}\sim 1/\sqrt{T_{\mathrm{obs}}}$
By only injecting BBHs signals into the time series and notching out the resolvable ones in the time-frequency domain as we did in Ref. Zhong et al. (2023, 2024), we find that the residual BBH foreground is $\sim 100$ times lower than the baseline noise $\sigma_{n}(f)$ at every frequency, even for 10 yrs observation, which justifies neglecting its contribution to the joint likelihood in Eq. (4). The uncertainty on the BNS foreground differs by about 17% between the inference with known and unknown SFR, and it is well below the baseline noise at all frequencies, for both known and unknown SFR, and for all observation times. This is consistent with the results discussed in the previous section: the BNS foreground is well constrained in both our inferences, allowing us to decouple it from the CGWB. By applying the $\sim 1/\sqrt{T_{\mathrm{obs}}}$ scaling on the CGWB uncertainty, we can project sensitivity estimates for longer observations than 10 days. For the flat CGWB, we find that we can exclude $\Omega_{\mathrm{ref}}=0$ at $5\,\sigma$ level for $\Omega_{\mathrm{ref}}=2.7\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$ ($2.5\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$) in the unknown (known) SFR case, where $T_{\mathrm{obs}}$ is the observation time in years.
@@
Acknowledgements We thank Will M. Farr, Thomas Callister, and Francesco Iacovelli for helpful discussions. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation (NSF) Grants PHY-0757058 and PHY-0823459. B.Z. is supported by the Fermi Research Alliance, LLC, acting under Contract No. DE-AC02-07CH11359. E.B. and L.R. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. H.Z. and V.M. are in part supported by the NSF grant PHY-2409173. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.

### v2

# Two-Step Procedure to Detect Cosmological Gravitational Wave Backgrounds with Next-Generation Terrestrial Gravitational-Wave Detectors

@@ Abstract
Several methods have been proposed to bypass the CBC foreground. Given the high fraction of individually resolved CBC signals with XG detectors, one could simply fit and subtract them from the data in the frequency domain to lower the foreground Regimbau et al. (2017); Sachdev et al. (2020). However, the recovery of detected signals is never perfect, and the pile-up of residuals from imperfect subtraction can produce an effective foreground that is comparable to the original SGWB before removal Zhou et al. (2023, 2022); Song et al. (2024). Techniques have been proposed to further reduce this effective foreground in the frequency domain, from applying a projection method Cutler and Harms (2006); Sharma and Harms (2020) to estimating and subtracting the expected residual power Pan and Yang (2023). Ref. Belgacem et al. (2024) showed that subtraction results in the frequency domain are actually significantly more optimistic if one abandons the linear-signal approximation in favor of more realistic estimates of the residuals. Moreover, the authors derive an optimal filter specifically targeted to the cosmological background search. They find that applying this match filtering procedure significantly lowers the impact of the CBC foreground in the search for a cosmological background, introducing a dependence on the specific spectral shape of such background. On the other hand, Refs. Zhong et al. (2023, 2024) proposed a time-frequency approach to notch-out the detected signals by masking the associated pixels. They showed that this method could reduce the CBC foreground by one order of magnitude even when residuals from imperfect recovery are taken into account Zhong et al. (2024).
@@
Astrophysical population and setup. For both BBHs and BNSs, we assume the same population models as in Refs. Zhong et al. (2023, 2024). We employ the PowerLaw+Peak model from the latest LIGO-Virgo-KAGRA (LVK) catalog Abbott et al. (2023) to characterize the BBH mass distribution, while we assume that the BNS masses are uniformly distributed within the range $1,2]\,\rm{M_{\odot}}$ Abbott et al. (2023); Landry and Read (2021).We assume zero spins and an isotropic distribution of the orbital orientation and sky position for all the binaries. To obtain the redshift distribution, we convolve the star formation rate (SFRFinkel et al. (2022)
@@
with a power-law time delay distribution $p(t_{d})\propto t_{d}^{p}$ with p=1p=-1. The source-frame merger rate $R_{m}(z)$ then reads
@@
Let us consider a set of data segments $\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}$ from $N_{\mathrm{obs}}$ individual BNS detections and the cross-correlation spectrum $\hat{C}^{\star}(f)$ after notching out the resolvable BBH signals (see the Supplement not for the definition of $\hat{C}^{\star}(f)$). Their joint likelihood can be written as Callister et al. (2020)
$\mathscr{L}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}},\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})=\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})\times\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})\,$, (4)
where $\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})$ is the usual hierarchical likelihood for population inference on the detected events Mandel et al. (2019); Taylor and Gerosa (2018), while $\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})$ is the Gaussian likelihood for stochastic searches Mandic et al. (2012); Callister et al. (2017). Here $\bm{\Lambda}\bm{}\mathrm{CGWB}=\{\Omega_{\mathrm{ref}},\alpha\}$ encodes the parameters characterizing the CGWB, while $\bm{\Lambda}\bm{}\mathrm{BNS}$ is the set of hyperparameters defining the BNS population, which affect both the individually detected events and the SGWB from unresolved BNS signals. We note that the above joint likelihood might not be valid in general for two reasons: (1) it ignores the BBH notching residual, which is a foreground; and (2) we use the BNS data in both the individual event and in the stochastic search, which could result in a bias. Regarding the first concern, since we notch the resolvable BBH events instead of subtracting them, the notching residual is rather weak (see the solid gray line in Fig. 3, and also Ref. Zhong et al. (2024)). To address the second concern, in Section IV of the Supplemental Material we perform a test to show that the bias is negligible. To simplify the inference, we assume that the BNS mass distribution is perfectly known and that the angular parameters are isotropically distributed. Hence, the only BNS hyperparameters $\bm{\Lambda}\bm{}\mathrm{BNS}$ that we are inferring are the ones characterizing the redshift distribution. In particular, we consider two cases:
@@
    A simplistic known SFR scenario where we assume that the SFR is completely known, and all that is left to determine are the local merger rate, $\mathcal{R}_{0}$, and the power-law index of the time-delay distribution, pp. Then the joint inference involves a total of four parameters $\{\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB}\}=\{\mathcal{R}_{0},p,\Omega_{\mathrm{ref}},\alpha\}$.
@@
In vanilla searches for a single, dominant SGWB, its SNR is straightforward to compute, and it is a good figure of merit to estimate whether the target SGWB is strong enough to be detectable relative to the detector noise Abbott et al. (2023); Allen and Romano (1999); Abbott et al. (2021a). In our case, instead, we are jointly fitting for the BNS hyperparameters from individually detected sources plus two distinct SGWBs, and our target is, in fact, the subdominant CGWB. Hence, we turn to a full Bayesian approach and we study the inferred posterior to find evidence for CGWB detection.
@@
Joint search: Known SFR $319^{+25}_{-24}$ $1.00^{+0.04}_{-0.04} -$ - - $1.92^{+0.58}_{-0.52}\times 10^{-11}$ $0.10^{+0.32}_{-0.37}$
Joint search: Unknown SFR $305^{+48}_{-47}$ - $2.24^{+0.19}_{-0.16}$ $1.43^{+0.28}_{-0.24}$ $1.64^{+0.07}_{-0.07}$ $2.04^{+0.66}_{-0.66}\times 10^{-11}$ $0.06^{+0.31}_{-0.39}$
Ideal search - - - - - $1.95^{+0.35}_{-0.34}\times 10^{-11}$ $0.14^{+0.29}_{-0.31}$
@@
Sensitivity of the CGWB search. In order to further assess the constraining power of our procedure, in Fig. 3 we compare the noise level for stochastic searches, the uncertainty on the BNS foreground, and the residual BBH foreground after notching out the resolved signals. The baseline noise level $\sigma_{n}(f)$ is computed through Eqs. (S1) and (S2) of the Supplement. The uncertainty on the inferred BNS foreground is estimated by computing energy spectra $\Omega_{\mathrm{BNS}}(f)$ for all posterior samples and the resulting standard deviation $\sigma_{\Omega_{\mathrm{BNS}}}(f)$ at each frequency (see the Supplement for the description of $\Omega_{\mathrm{BNS}}(f)$ calculation). Increasing the observation time leads to a larger number of individually observed BNS systems, resulting in a better estimate of the BNS population parameters and of $\sigma_{\Omega_{\mathrm{BNS}}}(f)$. By running our full simulation using fewer data, we explicitly verified that the width of the posterior distributions of BNS population parameters, as well as $\sigma_{\Omega_{\mathrm{BNS}}}$, scale as $1/\sqrt{T_{\mathrm{obs}}}$: $\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T^{\prime}_{obs}}}(f)=\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T_{obs}}}(f)\left(\sqrt{T_{\mathrm{obs}}}/\sqrt{T^{\prime}_{\mathrm{obs}}}\right)$. The sensitivity of stochastic searches also scales as $\sim 1/\sqrt{T_{\mathrm{obs}}}$ Allen and Romano (1999). We can, therefore, extrapolate our results from the 10 day simulation to much longer observation times, up to 10 years in Fig. 3, provided that the BBH notching residual remains subdominant.
By only injecting BBHs signals into the time series and notching out the resolvable ones in the time-frequency domain as we did in Ref. Zhong et al. (2023, 2024), we find that the residual BBH foreground is $\sim 100$ times lower than the baseline noise $\sigma_{n}(f)$ at every frequency, even for 10 yrs observation, which justifies neglecting its contribution to the joint likelihood in Eq. (4). To confirm that this residual foreground is indeed negligible when the observation time is much longer than 10 days, we also show the BBH notching residual considering a one-year-long observation in Fig. S1, and again find it much weaker than the detector sensitivity; hence, it can be neglected. The uncertainty on the BNS foreground differs by about 17% between the inference with known and unknown SFR, and it is well below the baseline noise at all frequencies, for both known and unknown SFR, and for all observation times. This is consistent with the results discussed in the previous section: the BNS foreground is well constrained in both our inferences, allowing us to decouple it from the CGWB. By applying the $\sim 1/\sqrt{T_{\mathrm{obs}}}$ scaling on the CGWB uncertainty, we can project sensitivity estimates for longer observations than 10 days. For the flat CGWB, we find that we can exclude $\Omega_{\mathrm{ref}}=0$ at $5\,\sigma$ level for $\Omega_{\mathrm{ref}}=2.7\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$ ($2.5\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$) in the unknown (known) SFR case, where $T_{\mathrm{obs}}$ is the observation time in years.
@@
Acknowledgements We thank Will M. Farr, Thomas Callister, and Francesco Iacovelli for helpful discussions. We also thank Joseph D. Romano for providing useful comments and suggestions on a first draft of this manuscript. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation (NSF) Grants PHY-0757058 and PHY-0823459. B.Z. is supported by the Fermi Research Alliance, LLC, acting under Contract No. DE-AC02-07CH11359. E.B. and L.R. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. H.Z. and V.M. are in part supported by the NSF grant PHY-2409173. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.